OpenStudy (anonymous):

find the limit: (n/(n+1))^n

OpenStudy (jamesj):

Hopefully you know this result \[ \lim_{n \rightarrow \infty} (1 + a/n)^n = e^a \] Given that, you can transform your question into something of this form.

OpenStudy (anonymous):

so my solution manual states that the answer is 1/e^5

OpenStudy (anonymous):

how would i go about solving for that?

OpenStudy (jamesj):

I don't agree that the answer is 1/e^5. Double check that or resign yourself to it being wrong. This limit is equal to 1/e

OpenStudy (anonymous):

wait my bad its supposed to be n+5* not n+1